OpenStudy (anonymous):

x - 3/4 = x/2 please solve the equation. thanks!

OpenStudy (anonymous):

x - 3/4 = x/2 4x - 3 = 2x 2x - 3 = 0 2x = 3 x = 3/2

OpenStudy (anonymous):

im sorry i meant to put... x-3 over 4 =x over 2